cocnsider the combination of two capacitors C1 and C2 with C2>C1 when connected in parallel , the equivalent capacitance is 8 times the equivalent capacitance of the same connected in series. calculate the ratio of capacitors, C2/C1.

1
Expert's answer
2021-12-15T16:21:59-0500

"C_1+C_2=8\frac{C_1C_2}{C_1+C_2}\to(\frac{C_2}{C_1})^2-6\cdot(\frac{C_2}{C_1})+1=0\to"


"C_2\/C_1=5.83" . Answer